SwRI CAPS ICO DOY 005 page
 
CAPS Checkout Report for 1999 DOY 005
From the desk of CAPS PI: Dave Young
Today's activities centered on bringing up the IMS sensor high voltages part way. This activity was entirely successful. The data push from JPL via the Internet was generally satisfactory throughout the pass. CAPS activities were shared during the pass with INMS and MIMI without any apparent difficulties.
The IBS and ELS sensors were left on after completion of DOY 004 activities. After playback of IBS and ELS data stored on the SSR, the two sensors were turned off preparatory to IMS activities. The two IMS microchannel plate (MCP) high voltage supplies were turned on first. Detector noise rates were very similar to those observed in the lab and followed the usual pattern of slow decay with time after initial switch-on. As with the IBS and ELS detectors, no noise has been observed in the detectors other than the expected random background component. This indicates that the s/c electromagnetic environment is very quiet in the present configuration. The IMS electrostatic analyzer (ESA) high voltage supply was then brought up and allowed to scan over its full range of 7500 volts. The MCP and ESA supplies were then turned off and the time-of-flight (TOF) acceleration and retardation high voltage supplies were turned on. These two supplies were brought up in parallel to +/- 9000 volts. (On the next pass the TOF supplies will be taken up to the full operating voltage of +/- 14,600 volts.) The MCP and ESA supplies were brought back up to their full operating voltages. IMS was then operational at the reduced TOF voltage. As expected, few (if any) ions were seen in the sensor because of the low TOF voltage and the orientation of the s/c away from the solar wind (see CAPS DOY 004 report). IBS and ELS were brought back into operation.
At this point the actuator was started again with a scan range of +/- 20 degrees in order to evaluate its performance following the DOY 004 "sticking" problem at -98 deg. Performance at +/- 20 deg. of actuation was acceptable so +/- 40 deg. was commanded on. The actuator was left in this mode for the remainder of the pass and the next day. Because of the sticking we are considering limiting the actuator scan range during the CAPS/MIMI roll into the solar wind. The IMS A-cycle timing issue that results in loss of 50% of the IMS data remains a problem. It is currently under investigation by GSFC. The DOY 005 data show variations in the ELS electron spectra that are correlated with the position of the CAPS actuator. The variations are caused by structure in the photoelectron dominated plasma sheath surrounding the Cassini spacecraft.
 
Back to ICO status page
 
Snapshot of Data from DOY 005, UTC: 15:32:42
Snapshot DOY 005 Data
ELS Data from DOY 005, UTC: 15:34:50
Snapshot DOY 005 Data